OpenStudy (mellamokatie):

For the following system of equations, what is the x- value of the solution? -x + 2y = 6 6y = x + 18 A. -1 B. 0 C. 1 D. 6

OpenStudy (mellamokatie):

@plainntall

OpenStudy (mellamokatie):

@sweetburger

OpenStudy (mellamokatie):

@Mehek14

Mehek (mehek14):

Do you want to use substitution or elimination

OpenStudy (mellamokatie):

What do you mean?

Mehek (mehek14):

To solve

OpenStudy (mellamokatie):

I dont know what that is sorry D:

Mehek (mehek14):

That's why I'm asking so I can show you how to use it

OpenStudy (mellamokatie):

oh ok

Mehek (mehek14):

So which one do you want to use?

OpenStudy (mellamokatie):

wait I think I know what they mean so substitution is where you pull answer's like A B C or D right? Lets do that if that's the case.

Mehek (mehek14):

No

Mehek (mehek14):

Substitution is when we use the value of either x or y from one equation and plug it into the other equation

OpenStudy (mellamokatie):

uhm okay

OpenStudy (mellamokatie):

so then that?

Mehek (mehek14):

Well are we going to use substitution or elimination?

OpenStudy (mellamokatie):

subsitution

Mehek (mehek14):

Solve for x in the second equation 6y = x + 18

OpenStudy (mellamokatie):

okay how?

Mehek (mehek14):

isolate x

OpenStudy (mellamokatie):

oh okay 18-6 = 12 how would that connect with the equation?

Mehek (mehek14):

I have no idea how you got 18 - 6 = 12

OpenStudy (mellamokatie):

ugh I don't know either then again I don't even know how to do this please help :/

Mehek (mehek14):

Ok to isolate x 6y = x + 18 you would subtract 18 on both sides 6y - 18 = x

OpenStudy (mellamokatie):

Thats what I was trying to do so I got 12

Mehek (mehek14):

you cannot subtract 6y - 18 and get 12

Mehek (mehek14):

all we were doing was solving for x

OpenStudy (mellamokatie):

okay so whats the answer

Mehek (mehek14):

Well we just isolated x so it's 6y - 18 = x now we substitute `6y - 18` instead of x in the first equation

OpenStudy (mellamokatie):

ok

Mehek (mehek14):

So substitute that into the first equation

OpenStudy (mellamokatie):

how would I do that? What does it look like?

Mehek (mehek14):

instead of x in the first equation, put 6y - 18

OpenStudy (mellamokatie):

so then 6y -18 +2y =6?

Mehek (mehek14):

-(6y - 18) + 2y = 6 -6y + 18 + 2y = 6 combine -6y + 2y =

OpenStudy (mellamokatie):

4?

Mehek (mehek14):

-4y

Mehek (mehek14):

So 18 - 4y = 6 6 - 18 = ?

OpenStudy (mellamokatie):

12

Mehek (mehek14):

-12

OpenStudy (mellamokatie):

oh ok

Mehek (mehek14):

-4y = -12 -12/-4 = ?

OpenStudy (mellamokatie):

-3

Mehek (mehek14):

3

Mehek (mehek14):

so y = 3

Mehek (mehek14):

put -3 instead of y in any equation and solve for x

OpenStudy (mellamokatie):

okay one moment

OpenStudy (mellamokatie):

-4y = -12 this equation?

Mehek (mehek14):

no

Mehek (mehek14):

the 2 you have

OpenStudy (mellamokatie):

oh so then -3 + 2y = 6

Mehek (mehek14):

-3 instead of y not x

OpenStudy (mellamokatie):

oh so then -x + 2(-3)= 6

OpenStudy (mellamokatie):

6(-3) = x + 18

Mehek (mehek14):

You just have to use one of them solve and you get x = 0

OpenStudy (mellamokatie):

Okay so the overall answer is 0?

Mehek (mehek14):

Yes

OpenStudy (mellamokatie):

Thank you I really appreteate that!!!

Mehek (mehek14):

You're welcome :)